OpenStudy (anonymous):

[6.03] Solve the following system. x − y = 10 x + y = 4 (−3, 7) (−7, 3) (7, 3) (7, −3)

OpenStudy (anonymous):

add bouth equations and see what you get

OpenStudy (callisto):

x − y = 10 -(1) x + y = 4 -(2) (1) + (2): (x − y) + (x + y) = 10 + 4 What do you get from here?

OpenStudy (anonymous):

Another option is to solve either equation for either y or x, then plug the result back into the equation using substitution.

OpenStudy (anonymous):

you get 14 ?

OpenStudy (anonymous):

x − y = 10 x + y = 4 ______________ + 2x = 14

OpenStudy (anonymous):

you can find value of x from here. Later use it to find value of y

OpenStudy (anonymous):

7

OpenStudy (anonymous):

right x=7, now plug this in any original equation and solve it for y

OpenStudy (anonymous):

Its 3 ?

OpenStudy (anonymous):

be carefull with signs

OpenStudy (anonymous):

7,3 or 7,-3

OpenStudy (anonymous):

if foe example you use this equation to find y: x + y = 4 you get: 7+y=4 so y=4-7=-3 and the answer will be: (7,-3)